### 30 Mar 2026 v2

# These $\equalcontThese authors contributed equally to this work.$

Abstract
In this note, we establish several results concerning the continuity (or weak convergence) of the complex Monge-Ampère operator on compact Hermitian manifolds. At the end of this note, we find a weak solution of the complex Monge-Ampère equation on a compact Hermitian manifold under the assumption of the existence of a smooth subsolution.
keywords:
$\omega$-plurisubharmonic functions, weakly convergent, the complex Monge-Ampère operator, compact Hermitian manifolds, convergence in capacity, weak solutions of the complex Monge-Ampère equation.
pacs:
MSC Classification]32U05, 32Q15, 32W20.
1 Introduction
1.1 Background
Let $(X,\omega)$ be a compact Hermitian manifold of complex dimension $n$, equipped with the Hermitian fundamental form $\omega$ given in local coordinates by $\omega=\frac{i}{2}\sum\limits_{\alpha,\beta}a_{\alpha,\bar{\beta}}dz_{\alpha}\wedge d\bar{z}_{\beta},$ where $(a_{\alpha,\bar{\beta}})$ is a positive definite Hermitian matrix. In the case $d\omega=0$, then $\omega$ is called a Kähler form and $(X,\omega)$ is said to be a compact Kähler manifold. The smooth volume form $dV$ associated to this Hermitian metric is given by the $n$-th wedge product $dV=\omega^{n}$. The study of the continuity of the complex Monge-Ampère operator $(dd^{c}\bullet)^{n}$ plays an important role in many problems of pluripotential theory. The original results in the direction of this research from the seminal work of Betford and Taylor in [1]. In this paper, Bedford and Taylor proved that if a sequence of locally bounded plurisubharmonic functions $(u_{j})\subset PSH(\Omega)\cap L^{\infty}_{loc}(\Omega)$, where $PSH(\Omega)$ denotes the set of plurisubharmonic functions on an open subset $\Omega\subset\mathbb{C}^{n}$, is decreasing to a locally bounded plurisubharmonic function $u$ then $(dd^{c}u_{j})^{n}$ is weakly convergent to $(dd^{c}u)^{n}$. They also established a result about weak convergence of the sequence of the complex Monge-Ampère $(dd^{c}u_{j})^{n}$ to $(dd^{c}u)^{n}$ under the assumption that $(u_{j})$ is an increasing sequence and a.e. convergent concerning the Lebesgue measure $dV_{2n}$ to a locally bounded plurisubharmonic $u$. In summary, by relying on the results obtained by Bedford-Taylor, we observe that the Monge-Ampère operator $(dd^{c}\bullet)^{n}$ is continuous in weak*-topology along monotonic sequences in the class of locally bounded plurisubharmonic functions.
It is worth noting, however, that monotone convergence of a sequence of plurisubharmonic functions $(u_{j})$ to a psh function $(u)$ does not always occur. This raises the natural question of whether the monotonicity assumption can be replaced by a weaker condition, such as convergence in $L^{p}$, which still guarantees the convergence in the weak*-topology of the sequence of complex Monge-Ampère operators $(dd^{c}u_{j})^{n}$ to $(dd^{c}u)^{n}$. Unfortunately, that does not happen. Due to a counterexample given in [3], Cegrell showed that there exists a sequence of continuous plurisubharmonic functions $(u_{j})$ in $\mathbb{C}^{2}$ converging in the $L^{p},1\leq p<+\infty$-topology to a plurisubharmonic function $u$, but the corresponding sequence of Monge-Ampère operators $(dd^{c}u_{j})^{2}$ does not converge in the weak*-topology to $(dd^{c}u)^{2}$. Thus, convergence in the $L^{p}$-topology for $p\geq 1$ does not imply the weak continuity of the sequence of Monge-Ampère operators.
At this point, one seeks a condition weaker than the monotone convergence of sequences of plurisubharmonic functions, yet strong enough to establish the continuity of the Monge-Ampère operator sequence in the weak*-topology. In this direction, Xing, in his 1996 paper [19], introduced the notion of convergence in $C_{n}$-capacity for sequences of plurisubharmonic functions. He proved that if $(u_{j})$ is a sequence of plurisubharmonic functions, uniformly bounded on an open subset $\Omega\subset\mathbb{C}^{n}$, and converging to a locally bounded plurisubharmonic function $u$ in $C_{n}$-capacity, then the corresponding Monge-Ampère measures $(dd^{c}u_{j})^{n}$ converge to $(dd^{c}u)^{n}$ in the weak-topology. Also in that paper, based on the concept of convergence in $C_{n}$-capacity, Xing obtained several other forms of weak convergence for sequences of Monge-Ampère measures. For more details on these results, we refer readers to [19]. Next, in [20], relying on the concept of convergence in $C_{n}$-capacity, Xing has provided deeper results concerning the relationship between the convergence in $C_{n}$-capacity of a sequence of uniformly bounded plurisubharmonic functions $(u_{j})$ to a locally bounded plurisubharmonic function $u$, and the weak convergence of the sequence of corresponding Monge-Ampère measures $(dd^{c}u_{j})^{n}$ and $(dd^{c}u)^{n}$. In fact, Xing found a type of convergence of the complex Monge-Ampère measures which is essentially equivalent to convergence in $C_{n}$-capacity of plurisubharmonic functions, respectively (see Theorems 2.8, 2.11, 2.14 and Corollaries 2.10, 2.12 in [20]).
Note that, from the monotone convergence of a locally bounded sequence of plurisubharmonic functions, it is not difficult to deduce the convergence of this sequence in $C_{n}$-capacity. This shows that convergence in $C_{n}$-capacity is a weaker condition than monotone convergence, yet it is still effective for establishing the weak convergence of Monge-Ampère measure sequences. It should be noted that it was precisely the discovery of the weak convergence of the Monge-Ampère measures of bounded plurisubharmonic functions on hyperconvex domains $\Omega\subset\mathbb{C}^{n}$, made between 1998 and 2004, that led Cegrell in [4] and [6] to show that the Monge-Ampère operator $(dd^{c}\bullet)^{n}$ is well-defined on certain classes of unbounded plurisubharmonic functions. He introduced and investigated some classes of plurisubharmonic functions such as $\mathcal{E}_{p}(\Omega),\mathcal{F}_{p}(\Omega),\mathcal{E}(\Omega),\mathcal{F}(\Omega)$, and, at the same time, he investigated the solvability of the Monge-Ampère equations within these classes. In [6], Cegrell showed that if $\Omega\subset\mathbb{C}^{n}$ is a hyperconvex domain, then $\mathcal{E}(\Omega)$ is the biggest class on which the Monge-Ampère operator is well-defined and continuous under decreasing sequences of plurisubharmonic functions in that class. For more details on the aforementioned issues, we refer readers to the concepts and results presented in [4] and [6], respectively.
A natural question that arises here is whether the convergence in $C_{n}$-capacity of plurisubharmonic functions in the above-mentioned Cegrell’s classes implies the weak convergence of the corresponding Monge-Ampère measures $(dd^{c}\bullet)^{n}$. The study of the weak convergence of the complex Monge-Ampère operator in these classes was initiated by Cegrell in a 2001 manuscript (see [5]) and was completed and published in [7] in 2012. In [7], he proved that if $(u_{j})$ is a sequence of plurisubharmonic functions in $\mathcal{E}(\Omega)$, where $\Omega$ is a hyperconvex domain in $\mathbb{C}^{n}$, $u\in\mathcal{E}(\Omega)$ and there exists $u_{0}\in\mathcal{E}(\Omega)$ such that $u_{j}\geq u_{0}$ for all $j\geq 1$, $(u_{j})$ is convergent to $u$ in $C_{n}$-capacity then $(dd^{c}u_{j})^{n}$ is weakly convergent to $(dd^{c}u)^{n}$. Since then, the problem of studying the weak convergence of the Monge-Ampère operator sequence in Cegrell classes under the influence of convergence in $C_{n}$-capacity has attracted the attention of several authors.
In [20], Xing investigated the weak convergence of a sequence of Monge-Ampère operator $(dd^{c}\bullet)^{n}$ in the class $\mathcal{F}(\Omega)$. Pham Hoang Hiep in [13], by removing the condition $u_{j}\geq u_{0}$, it was shown in [7] that if $(u_{j})\subset\mathcal{E}(\Omega)$ and $u_{j}\rightarrow u$ in $C_{n}$-capacity, then
$\lim\inf\limits_{j\to\infty}(dd^{c}u_{j})^{n}\geq 1_{\{u>-\infty\}}(dd^{c}u)^{n}.$
The results discussed above concerning convergence in $C_{n}$-capacity and the associated weak convergence of Monge–Ampère measures have thus far been formulated in the local setting, i.e., on open subsets of $\mathbb{C}^{n}$.
In this work, we aim to extend this line of research to the global setting, with a particular focus on the connection between convergence in $C_{n}$-capacity and the weak convergence of Monge–Ampère operators on complex manifolds, specifically, on compact Kähler manifolds and, more generally, on compact Hermitian manifolds. These developments are relatively recent and serve as the principal motivation for the present paper. We begin by investigating the weak convergence of Monge–Ampère measures under convergence in capacity on a compact Kähler manifold $(X,\omega)$ of complex dimension $n$. Since every plurisubharmonic function on $X$ is necessarily constant, the appropriate framework in this context is to work with quasi-plurisubharmonic functions, also known as $\omega$-plurisubharmonic functions (abbreviated as $\omega$-psh functions). This class of functions was introduced and studied in detail, and some fundamental results for this class were obtained in [10]. By $PSH(\omega)$ we denote the class of $\omega$-psh on $(X,\omega)$. We recall some main results in [10]. If $u\in PSH(\omega)\cap L^{\infty}(X)$ then in[10] they showed that there exists a Monge-Ampère measure $\omega_{u}^{n}=(\omega+dd^{c}u)^{n}$ and it is continuous on decreasing sequences ( also see [15] for arguments on the existence of Monge-Ampère measure $\omega_{u}^{n}$ in the situation $(X,\omega)$ is a compact Hermitian manifold). Definition 3.4 in [10] deals with the notion of capacity of a Borel subset $E\subset X$ on a compact Kähler manifold $(X,\omega)$. Corollary 3.8 gives the quasi-continuity of $\omega$-psh. Remark that here we only recall some notions concerning the class of $\omega$-psh functions. The precise definition of $\omega$-plurisubharmonic functions as well as the notion of capacity in the case where $(X,\omega)$ is a compact Hermitian manifold will be presented in detail in the following section of the paper.
Next, in [11], Guedj and Zeriahi introduced and studied the class $\mathcal{E}(X,\omega)$. They showed that the complex Monge-Ampère operator $(\omega+dd^{c}\bullet)^{n}$ is well defined on the class $\mathcal{E}(X,\omega)$ of $\omega$-psh with finite weighted Monge-Ampère energy. The class $\mathcal{E}(X,\omega)$ is the largest class of $\omega$-plurisubharmonic ($\omega$-psh) functions on which the complex Monge–Ampère operator is well defined and for which the comparison principle holds. For detailed results concerning this class, we refer the reader to [11]. The study of convergence with respect to the global capacity $Cap_{\omega}$ on compact Kähler manifolds was initiated in [12, 14]. In that paper, Hiep established several equivalent characterizations for convergence in $Cap_{\omega}$ of a uniformly bounded sequence $(u_{j})$ of $\omega$-psh functions to a given $\omega$-psh function $u$ (see Theorem 2.1 and Lemma 2.3 in [12]).
Subsequently, the study of the weak continuity of the Monge-Ampère operator $(\omega+dd^{c}\bullet)^{n}$ concerning convergence in $Cap_{\omega}$ on the class $\mathcal{E}(X,\omega)$ was carried out by Xing in [20]. Up until 2008, the results obtained by Xing in [20, 21] represented the most profound and significant contributions to this line of research. Subsequently, in 2012, Dinew and Hiep further advanced the study of convergence in capacity on compact Kähler manifolds within the class $\mathcal{E}(X,\omega)$ (see [8]). Their work synthesizes and extends several results previously established in [20]. The results mentioned above are results concerning the continuity of the Monge-Ampère operator for $\omega$-plurisubharmonic functions related to convergence in the capacity $Cap_{\omega}$ on compact Kähler manifolds.
Note that, when studying the continuity of the Monge–Ampère operator $(dd^{c}u_{j})^{n}$ on Cegrell classes $\mathcal{E}(\Omega)$, it is usually necessary to assume that the sequence of functions $(u_{j})$ is bounded from below by a function $u_{0}$. However, when considering the continuity problem for the Monge–Ampère operator $(\omega+dd^{c}\bullet)^{n}$ on compact Kähler manifolds, as in the papers of Xing (refer [20, 21]) or Dinew and Hiep in [8], the authors were able to dispense with this condition, since finding such a lower bounding function is not always feasible. Now, regarding compact Hermitian manifolds, what results have been obtained in this direction of research? Unfortunately, results in this direction for the case $(X,\omega)$, which is a compact Hermitian manifold, remain quite limited up to now. Only recently one knew the study of weak convergence of the Monge-Ampère operator for the class of $\omega$-plurisubharmonic functions appeared, notably in the paper by Kołodziej and Ngoc Cuong Nguyen in [17]. Regrettably, in this paper, under the assumptions of convergence in capacity, the authors were only able to obtain weak convergence of a subsequence $(\omega^{n}_{u_{j_{k}}})$ of the original sequence $(\omega^{n}_{u_{j}})$ (see Theorem 1.1 in [17]). Inspired by the results in [17], in this paper, we aim to provide additional results on the weak convergence of the Monge-Ampère operator on compact Hermitian manifolds for the class of $\omega$-plurisubharmonic functions, in relation to convergence in $Cap_{\omega}$.
1.2 Organization
The paper is presented in four sections and organized as follows.
In Section 2, we present important results concerning the class of $\omega$-plurisubharmonic functions ($\omega$-psh) that will be used in the subsequent parts of the paper. This includes the definition of the $\omega$-psh class, the Monge-Ampère operator for bounded $\omega$-psh functions, the notion of weak convergence of a sequence of Monge-Ampère operators, the definition of the capacity $Cap_{\omega}$ and the convergence of a sequence $(u_{j})$ of $\omega$-psh functions to an $\omega$-psh function in $Cap_{\omega}$, as well as several other related concepts. Section 3 is devoted to investigating the weak convergence of a sequence of the complex Monge-Ampère operator $(\omega+dd^{c}u_{j})^{n}$ to $(\omega+dd^{c}u)^{n}$ if we assume that $u_{j}$ is convergent to $u$ in $Cap_{\omega}$. In Section 4, we establish the existence of weak solutions of the Monge-Ampère equations on compact Hermitian manifolds under the assumption that there exists a smooth subsolution.
Acknowledgments. The authors would like to thank the anonymous referees for their valuable comments and suggestions, which led to improvements in the paper’s exposition. The first and second named authors are supported by Grant number B2025-CTT-10 from the Ministry of Education and Training, Vietnam. This work was written during our visit to the Vietnam Institute for Advanced Study in Mathematics (VIASM) in the Spring of 2025. We also thank VIASM for financial support and hospitality.
2 Preliminairies
Through this paper by $(X,\omega)$ we denote a fixed compact Hermitian manifold $X$ of dimension $n$ equipped with a Hermitian metric $\omega$. As mentioned in the beginning of the paper, locally, $\omega$ can be written as $\omega=\frac{i}{2}\sum\limits_{\alpha,\beta}a_{\alpha,\bar{\beta}}dz_{\alpha}\wedge d\bar{z}_{\beta},$ where $(a_{\alpha,\bar{\beta}})$ is a positive definite Hermitian matrix. The smooth volume form $dV$ associated to this Hermitian metric is given by the $n$-th wedge product $dV=\omega^{n}$. Moreover, we assume that $\int\limits_{X}dV<\infty$.
Definition 2.1.
A function $u:X\rightarrow[-\infty,+\infty)$ is call $\omega$-plurisubharmonic function (briefly, $\omega$-psh) if it satisfies the following conditions:
(i) $u$ is upper semi-continuous on $X$.
(ii) $u\in L^{1}(X,dV)$ and $\omega+dd^{c}u\geq 0$ in the sense of current.
By $PSH(X,\omega)$ or $PSH(\omega)$ we denote the set of $\omega$-psh functions on $X$. Now we deal with the complex Monge-Ampère operator on $PSH(\omega)$. Let $u\in PSH(\omega)\cap L^{\infty}(X)$. Following the fundamental work of Bedford and Taylor in [1], we know that the complex Monge-Ampère operator $(\omega+dd^{c}u)^{n}$ is well-defined. Moreover, it is a positive Borel measure on $X$. For a detailed account of the existence of the operator $(\omega+dd^{c}u)^{n}$, where $u\in\text{PSH}(\omega)\cap L^{\infty}(X)$, we refer readers to the construction provided before Proposition 2.1 in [9] and more detail, in Proposition 1.2 of [15].
Definition 2.2.
Let $(u_{j})\subset PSH(\omega)\cap L^{\infty}(X)$ and $u\in PSH(\omega)\cap L^{\infty}(X)$. The sequence of Monge-Ampère operators $(\omega+dd^{c}u_{j})^{n}$ is called weak convergence to $(\omega+dd^{c}u)^{n}$ if for all $\psi\in C^{\infty}_{0}(X)$ the following condition holds
$\lim\limits_{j\to\infty}\int\limits_{X}\psi(\omega+dd^{c}u_{j})^{n}=\int\limits_{X}\psi(\omega+dd^{c}u)^{n}.$
where $C^{\infty}_{0}(X)$ is the set of smooth functions with compact support in $X$. Each $\psi\in C^{\infty}_{0}(X)$ also is called a test function. When the sequence of Monge-Ampère operators $(\omega+dd^{c}u_{j})^{n}$ is convergent weakly to $(\omega+dd^{c}u)^{n}$, we also say that the Monge-Ampère operator is continuous, where “continuity ”is understood in the sense that it is continuous in the weak*-topology.
Now as in [10] we give the notion of capacity $Cap_{\omega}$ of a Borel subset $E\subset X$.
Definition 2.3.
Assume that $E$ is a Borel subset of $X$. Capacity of $E$ denoted by $Cap_{\omega}(E)$ and is defined by
$Cap_{\omega}(E)=\sup\Bigl\{\int\limits_{E}(\omega+dd^{c}u)^{n}:u\in PSH(\omega),\ \ 0\leq u\leq 1\Bigl\}.$
Next, we deal with the convergence in capacity of a sequence $(u_{j})\subset PSH(\omega)$ to a function $u\in PSH(\omega)$
Definition 2.4.
A sequence $(u_{j})$ of $\omega$-psh functions on $X$ is said to be convergence in capacity to a $\omega$-psh $u$ if for all $\eta>0$ the following condition holds
$\lim\limits_{j\to\infty}Cap_{\omega}\Bigl\{|u_{j}-u|>\eta\Bigl\}=0.$
By repeating the same arguments as in the proof of Corollary 3.8 in [10] and arguments are explained before Proposition 2.1 in [9], we note that the quasi-continuity of $\omega$-plurisubharmonic functions is still valid in Hermitian setting. Namely, we have the following result.
Proposition 2.5.
Let $u\in PSH(\omega)$ be a $\omega$-psh function. Then for each $\varepsilon>0$ there exists an open subset $G\subset X$ with $Cap_{\omega}(G)<\varepsilon$ and $u|_{X\setminus G}$ is continuous.
Next, we recall the notion of variation of a measure. Given a Borel measure $\mu$ (non-necessarily positive). One defines its variation $\|\mu\|$ on a Borel set $A$ by
$\|\mu\|(A)=\sup\Bigl\{\sum\limits_{j=1}^{\infty}|\mu(A_{j})|\Bigl\},$
where $\{A_{j}\},j\geq 1$ is any disjoint at most countable partition of the set $A$ by Borel sets. The classical Hahn decomposition theorem states that in fact that it is enough to consider only the special partition given by the two Hahn sets,intuitively,the pieces where $\mu$ is positive and negative. To avoid lengthy expressions, we adopt the convention that from now on, when writing the Monge-Ampère operator of a function $u\in PSH(\omega)\cap L^{\infty}(X)$, we use the notation $\omega^{n}_{u}$. Thus, we understand $\omega^{n}_{u}=(\omega+dd^{c}u)^{n}$.
3 Continuity of the complex Monge-Ammpère operator on compact Hermitian manifolds
First, we recall the following class of $\omega$-sh functions introduced and investigated in [17]. Namely, let $(X,\omega)$ be a compact Hermitian manifold mentioned in Section 2. By $\mathcal{P}_{0}$ we denote the set of $\omega$-psh functions defined by
$\mathcal{P}_{0}=\Bigl\{v\in PSH(\omega)\cap L^{\infty}(X):\ \ \sup\limits_{X}v=0\Bigl\}.$
We begin the section by proving a following result which serves as an essential tool for subsequent results.
Theorem 3.1.
Let $(u_{j})\subset PSH(\omega)\cap\mathcal{P}_{0}$ be a uniformly bounded sequence of $\omega$-psh. Assume that $(u_{j})$ is convergent to $u\in\mathcal{P}_{0}$ in $Cap_{\omega}$. Then $\omega^{n}_{u_{j}}$ is weakly convergent to $\omega^{n}_{u}$.
Proof.
It is enough to show that every subsequence of the $\omega^{n}_{u_{j}}$ has its subsequence which is weakly convergent to $\omega^{n}_{u}$. Therefore, without loss of generality, we may assume that $(u_{j})$ is a subsequence of the original sequence. From the hypothesis $(u_{j})$ is convergent in $Cap_{\omega}$ to $u$, by Proposition 2.5 in [17] we have
$\lim\limits_{j\to\infty}\int\limits_{X}|u_{j}-u|\omega^{n}_{u_{j}}=0.$ (3.1)
Hence, coupling (3.1) and Lemma 2.4 in [17] it follows that there exists a subsequence $(u_{j_{k}})$ of the $(u_{j})$ such that $\omega^{n}_{u_{j_{k}}}$ is weakly convergent to $\omega^{n}_{u}$ and the desired conclusion follows. ∎
Now we give the following result. Note that this result was mentioned in [10] and [9], but no proof was given there. By relying on the quasi-continuity of $\omega$-psh functions and Theorem 3.1, we will provide a proof here in Hermitian setting. Namely, we have
Corollary 3.2.
Let $(u_{j})$ be a uniformly bounded sequence of $\omega$-psh functions in $\mathcal{P}_{0}$. Assume that $(u_{j})$ is decreasing to $u\in\mathcal{P}_{0}$. Then $\omega^{n}_{u_{j}}$ is weakly convergent to $\omega^{n}_{u}$.
Proof.
By Theorem 3.1 it remains to show that $(u_{j})$ is convergent to $u$ in $Cap_{\omega}$. For $\varepsilon>0$, by the quasi-continuity of $u$ there exists an open subset $G\subset X$ such that $Cap_{\omega}(G)<\varepsilon$ and $u$ is continuous on $X\setminus G$. Note that $X\setminus G$ is compact, by Hartogs lemma for $\eta>0$ there exits $j_{0}>0$ such that $\forall\ \ j\geq j_{0}$ and for all $x\in X\setminus G$ we get that
$u_{j}(x)\leq u(x)+\eta.$ (3.2)
On the other hand, by the hypothesis we have
$u_{j}(x)\geq u(x)>u(x)-\eta,\forall\ \ x\in X.$
This fact together with (3.2) implies that for all $x\in X\setminus G$ and $j\geq j_{0}$ we deduce that
$u(x)-\eta<u_{j}(x)\leq u(x)+\eta.$ (3.3)
From (3.3) it follows that
$|u_{j}(x)-u(x)|<\eta,\forall\ \ j\geq j_{0},\ \ x\in X\setminus G.$
Hence, $\Bigl\{|u_{j}(x)-u(x)|>\eta\Bigl\}\subset G$ for $j\geq j_{0}.$ This yields
$Cap_{\omega}\Bigl(\Bigl\{|u_{j}(x)-u(x)|>\eta\Bigl\}\Bigl)\leq Cap_{\omega}(G)<\varepsilon,$
for $j\geq j_{0}$. Such as, the sequence $(u_{j})$ is convergent to $u$ in $Cap_{\omega}$ as we wanted. The proof is complete. ∎
The following result is an extension of Theorem 2 in [20] in Hermitian setting.
Theorem 3.3.
Assume that for all $v\in PSH(\omega)\cap L^{\infty}(X)$ the following condition holds
$\int\limits_{X}\omega^{n}_{v}=\int\limits_{X}\omega^{n}.$ (3.4)
Let $(u_{j})\subset\mathcal{P}_{0}$ be a uniformly bounded sequence of $\omega$-psh functions and $u\in\mathcal{P}_{0}$. Suppose that $u_{j}\rightarrow u$ in $L^{1}(X,dV)$ and for all $\delta>0$, $\lim\limits_{j\to\infty}\int\limits_{\{u\geq u_{j}+\delta\}}\omega^{n}_{u_{j}}=0$. Then $\omega^{n}_{u_{j}}\rightarrow\omega^{n}_{u}$ weakly.
Proof.
It is enough to show that any subsequence of the $(\omega^{n}_{u_{j}})$ has its subsequence which is weakly convergent to $\omega^{n}_{u}$. So we may assume that
$\int\limits_{\{u\geq u_{j}+\frac{1}{j}\}}\omega^{n}_{u_{j}}\rightarrow 0\ \text{as}\ j\to\infty.$
Put $\phi_{j}=\max\{u_{j},u-\frac{1}{j}\}$. Then $\phi_{j}\in\mathcal{P}_{0}$ and the sequence $(\phi_{j})$ is uniformly bounded on $X$. Now we prove that $\phi_{j}$ is convergent to $u$ in $Cap_{\omega}$. Indeed, we have
$\left|u_{j}-\left(u-\frac{1}{j}\right)\right|=\phi_{j}-u_{j}+\phi_{j}-\left(u-\frac{1}{j}\right).$ (3.5)
Since $u_{j}\rightarrow u$ in $L^{1}(X,dV)$ then $\int\limits_{X}|u_{j}-u|dV\rightarrow 0$ as $j\to\infty$. On the other hand, we have
$\int\limits_{X}\left|u_{j}-\left(u-\frac{1}{j}\right)\right|dV\leq\int\limits_{X}|u_{j}-u|dV+\frac{1}{j}\int\limits_{X}dV\rightarrow 0,\ \ \text{as $j\to\infty$.}$
However, by (3.5) it follows that
$0\leq\phi_{j}-\left(u-\frac{1}{j}\right)\leq\left|u_{j}-\left(u-\frac{1}{j}\right)\right|,$
then we get that $\phi_{j}\rightarrow u$ in $L^{1}(X,dV)$. By Proposition 2.5 for $\varepsilon>0$ there exists an open $G\subset X$ with $Cap_{\omega}(G)<\varepsilon$ and $u$ is continuous on $X\setminus G$. Using Hartogs lemma for $\eta>0$ we can find $j_{0}>0$ such that for $j\geq j_{0}$, $\frac{1}{j}<\eta$ and
$\phi_{j}\leq u+\eta,$
on $X\setminus G$. But we have $\phi_{j}\geq u-\frac{1}{j}>u-\eta$ on $X$. It follows that on $X\setminus G$ we get that
$|\phi_{j}-u|<\eta,\ \ \forall j\geq j_{0}.$
Hence, using the arguments in the proof of Corollary 3.2 we infer that $\phi_{j}\rightarrow u$ in $Cap_{\omega}$ as $j\to\infty$. Theorem 3.1 implies that $\omega^{n}_{\phi_{j}}\rightarrow\omega^{n}_{u}$ weakly in $X$ as $j\to\infty$. On the other hand, by the proof of Proposition 3.1 in [9] we get that
$1_{\{u_{j}>u-\frac{1}{j}\}}\omega^{n}_{\phi_{j}}=1_{\{u_{j}>u-\frac{1}{j}\}}\omega^{n}_{u_{j}}.$ (3.6)
From (3.6) we infer that
$\omega^{n}_{u_{j}}-\omega^{n}_{\phi_{j}}$ $=1_{\{u_{j}\leq u-\frac{1}{j}\}}\Bigl(\omega^{n}_{u_{j}}-\omega^{n}_{\phi_{j}}\Bigl)=-1_{\{u_{j}\leq u-\frac{1}{j}\}}\omega^{n}_{\phi_{j}}+1_{\{u_{j}\leq u-\frac{1}{j}\}}\omega^{n}_{u_{j}}.$
By the hypothesis
$\int\limits_{\{u_{j}\leq u-\frac{1}{j}\}}\omega^{n}_{u_{j}}=\int\limits_{\{u\geq u_{j}+\frac{1}{j}\}}\omega^{n}_{u_{j}}\rightarrow 0\ \text{as}\ j\to\infty.$
On the other hand, by using (3.4) and (3.6) we get that
$\int\limits_{X}1_{\{u_{j}\leq u-\frac{1}{j}\}}\omega^{n}_{\phi_{j}}$ $=\int\limits_{X}\omega^{n}_{\phi_{j}}-\int\limits_{\{u_{j}>u-\frac{1}{j}\}}\omega^{n}_{\phi_{j}}=\int\limits_{X}\omega^{n}-\int\limits_{\{u_{j}>u-\frac{1}{j}\}}\omega^{n}_{\phi_{j}}$
$=\int\limits_{X}\omega^{n}_{u_{j}}-\int\limits_{\{u_{j}>u-\frac{1}{j}\}}\omega^{n}_{u_{j}}=\int\limits_{\{u-\frac{1}{j}\geq u_{j}\}}\omega^{n}_{u_{j}}=\int\limits_{\{u\geq u_{j}+\frac{1}{j}\}}\omega^{n}_{u_{j}}\rightarrow 0,$
as $j$ tends to $\infty$ by the hypothesis. Hence, from the arguments presented above, we infer that $\omega^{n}_{u_{j}}-\omega^{n}_{\phi_{j}}\rightarrow 0$ weakly in $X$ as $j\to\infty$. However, we have
$\omega^{n}_{u_{j}}-\omega^{n}_{u}=\omega^{n}_{u_{j}}-\omega^{n}_{\phi_{j}}+\omega^{n}_{\phi_{j}}-\omega^{n}_{u}\rightarrow 0,$
weakly in $X$ as $j\to\infty$. Thus we achieve that $\omega^{n}_{u_{j}}\rightarrow\omega^{n}_{u}$ weakly as $j\to\infty$ as we wanted. The proof is complete. ∎
From Theorem 3.3 we obtain the following corollary which is an extension of Corollary 2 in [20] in Hermitian setting.
Corollary 3.4.
Let the condition (3.4) be satisfied and $(u_{j})\subset\mathcal{P}_{0}$ be a uniformly bounded sequence of $\omega$-psh functions and $u\in\mathcal{P}_{0}$, $u_{j}\rightarrow u$ in $L^{1}(X,dV)$. If $\omega^{n}_{u_{j}}=f_{j}dV$ with $\sup\limits_{j}\int\limits_{X}f_{j}^{p}dV<\infty,p>1$. Then $\omega^{n}_{u_{j}}$ is weakly convergent to $\omega^{n}_{u}$ as $j\to\infty$ in $X$.
Proof.
For each $\delta>0$ by using the Hölder inequality we get that
$\int\limits_{\{u\geq u_{j}+\delta\}}\omega^{n}_{u_{j}}=\int\limits_{\{u\geq u_{j}+\delta\}}f_{j}dV$ $\leq\Bigl(\int\limits_{X}f_{j}^{p}dV\Bigl)^{\frac{1}{p}}\Bigl(\int\limits_{\{u\geq u_{j}+\frac{1}{j}\}}dV\Bigl)^{1-\frac{1}{p}}$
$\leq\delta^{\frac{1}{p}-1}\sup\limits_{j}\Bigl(\int\limits_{X}f_{j}^{p}dV\Bigl)^{\frac{1}{p}}\Bigl(\int\limits_{X}|u_{j}-u|dV\Bigl)^{1-\frac{1}{p}}\rightarrow 0,$
as $j\to\infty$ by the hypothesis $u_{j}\to u$ in $L^{1}(X,dV)$. Hence, by Theorem 3.3 the proof is complete. ∎
Next, we have the following result.
Corollary 3.5.
Let the condition (3.4) be satisfied and $(u_{j})\subset\mathcal{P}_{0}$ be a uniformly bounded sequence of $\omega$-psh functions and $u\in\mathcal{P}_{0}$. Assume that $u_{j}\rightarrow u$ in $L^{1}(X,dV)$ and there exists $C>0$ such that for all $j\geq 1$, $\omega^{n}_{u_{j}}\leq CdV$. Then $\omega^{n}_{u_{j}}\rightarrow\omega^{n}_{u}$ weakly as $j\to\infty$ in $X$.
Proof.
Indeed, for $\delta>0$ we have
$\int\limits_{\{u\geq u_{j}+\delta\}}\omega^{n}_{u_{j}}\leq C\int\limits_{\{u\geq u_{j}+\delta\}}dV\leq\frac{C}{\delta}\int\limits_{X}|u-u_{j}|dV\rightarrow 0,$
as $j\to\infty$ and the desired conclusion follows. ∎
By using the same arguments as in the proof of Corollary 3.5 we note that Corollary 3.5 holds for a positive Radon measure $\mu$ on $X$. Namely, we have.
Corollary 3.6.
Let the condition (3.4) be satisfied and $(u_{j})\subset\mathcal{P}_{0}$ be an uniformly bounded sequence of $\omega$-psh functions and $u\in\mathcal{P}_{0}$. Assume $\mu$ is a positive Radon measure on $X$ with $\mu(X)<\infty$. If $u_{j}\rightarrow u$ in $L^{1}(X,d\mu)$ and $\omega^{n}_{u_{j}}\leq Cd\mu,C>0$ for all $j\geq 1$. Then $\omega^{n}_{u_{j}}$ is weakly convergent to $\omega^{n}_{u}$ in $X$.
Next, we have a notable result concerning the weak continuity of a sequence of Monge-Ampère operator in Hermitian setting.
Proposition 3.7.
Let the condition (3.4) be satisfied and $(u_{j})\subset\mathcal{P}_{0}$ be a uniformly bounded sequence of $\omega$-psh functions and $u\in\mathcal{P}_{0}$. Assume that $u_{j}\rightarrow u$ in $L^{1}(X,dV)$ and $\omega^{n}_{u_{j}}$ vanishes on the set $\{u_{j}<u\}$ for all $j$. Then $\omega^{n}_{u_{j}}$ is weakly convergent to $\omega^{n}_{u}$ in $X$.
Proof.
First, we prove $u_{j}\geq u$ on $X$ for all $j$. By the hypothesis $u\leq 0$ on $X$. Given $\varepsilon>0$. Choose $1>s>0$ such that $s\max\{1,\sup\limits_{X}|u|\}<\varepsilon$. Assume that $\varphi\in PSH(\omega)$ with $1\leq\varphi\leq 0$. Due to the condition (3.4), by Proposition 3.1 in [9], the comparison principle holds for bounded $\omega$-plurisubharmonic functions. That is, we have
$\int\limits_{\{u<v\}}\omega^{n}_{v}\leq\int\limits_{\{u<v\}}\omega^{n}_{u},$ (3.7)
for $u,v\in PSH(\omega)\cap L^{\infty}(X)$. Now for all $j\geq 1$, by using the comparison principle (3.7) we get that
$s^{n}\int\limits_{\{u_{j}<u-2\varepsilon\}}\omega^{n}_{\varphi}$ $\leq\int\limits_{\{u_{j}<u-2\varepsilon\}}(\omega+dd^{c}((1-s)u+s\varphi)))^{n}$
$\leq\int\limits_{\{u_{j}<(1-s)u+s\varphi-\varepsilon\}}\Bigl(\omega+dd^{c}\Bigl[(1-s)u+s\varphi\Bigl]\Bigl)^{n}$
$\leq\int\limits_{\{u_{j}<(1-s)u+s\varphi-\varepsilon\}}(\omega+dd^{c}u_{j})^{n}\leq\int\limits_{\{u_{j}<u\}}\omega^{n}_{u_{j}}=0.$
Taking the supremum over all such $\varphi$, by the definition of $Cap_{\omega}$, we get that
$Cap_{\omega}\Bigl(\{u_{j}<u-2\varepsilon\}\Bigl)=0\ \text{for all}\ j.$
Hence, $u_{j}\geq u-2\varepsilon$ on $X$ for all $j$. Let $\varepsilon\searrow 0$ we deduce that $u_{j}\geq u$ on $X$ for all $j$ and the desired conclusion follows. By repeating the same arguments as in the proof of Corollary 3.2 we infer that $u_{j}$ is convergent to $u$ in $Cap_{\omega}$. By Theorem 3.1, $\omega^{n}_{u_{j}}\rightarrow\omega^{n}_{u}$ weakly on $X$. The proof is complete. ∎
From the proof of Proposition 3.7 we also have the following result.
Corollary 3.8.
Let the condition (3.4) be satisfied and $(u_{j})\subset\mathcal{P}_{0}$ be an uniformly bounded sequence of $\omega$-psh functions and $u\in\mathcal{P}_{0}$. Assume that $u_{j}\rightarrow u$ in $L^{1}(X,dV)$ and $Cap_{\omega}\Bigl(\{u_{j}<u\}\Bigl)=0$ for all $j$. Then $\omega^{n}_{u_{j}}$ is weakly convergent to $\omega^{n}_{u}$ in $X$.
The following result is a form of Theorem 3.6 in [8] in Hermitian setting.
Proposition 3.9.
Let $(u_{j})\subset\mathcal{P}_{0}$ be an uniformly bounded sequence of $\omega$-psh functions and $u\in\mathcal{P}_{0}$. Assume that the following conditions hold:
(i) $u_{j}\rightarrow u$ in $L^{1}(X,dV)$
(ii) Let $\phi_{j}=\max\{u_{j},u\}$ and assume that
$\lim\limits_{j\to\infty}\int\limits_{X}\|\omega^{n}_{u_{j}}-\omega^{n}_{\phi_{j}}\|=0.$
Then $\omega^{n}_{u_{j}}\rightarrow\omega^{n}_{u}$ weakly.
Proof.
It is easy to see that $\phi_{j}\in\mathcal{P}_{0}$ and $(\phi_{j})$ is a uniformly bounded sequence of $\omega$-psh functions. From (i) we deduce
$\lim\limits_{j\to\infty}\int\limits_{X}|u_{j}-u|dV=0.$ (3.8)
Since $|u_{j}-u|=(\phi_{j}-u_{j})+(\phi_{j}-u)$ then it follows that $0\leq\phi_{j}-u\leq|u_{j}-u|$. Using (3.8), we infer that
$0\leq\int\limits_{X}(\phi_{j}-u)dV\leq\int\limits_{X}|u_{j}-u|dV\rightarrow 0,$
as $j\to\infty$. This yields that $\phi_{j}\rightarrow u$ in $L^{1}(X,dV)$. However, $\phi_{j}\geq u$ then by repeating the same arguments as in the proof of Corollary 3.2 we get that $\phi_{j}$ is convergent to $u$ in $Cap_{\omega}$. Theorem 3.1 implies that $\omega^{n}_{\phi_{j}}\rightarrow\omega^{n}_{u}$ weakly in $X$. We obtain
$\omega^{n}_{u_{j}}-\omega^{n}_{u}=\omega^{n}_{u_{j}}-\omega^{n}_{\phi_{j}}+\omega^{n}_{\phi_{j}}-\omega^{n}_{u}.$
Hence, it remains to prove that $\omega^{n}_{u_{j}}-\omega^{n}_{\phi_{j}}\rightarrow 0$ weakly in $X$. Let $\psi\in C^{\infty}_{0}(X)$ be an arbitrary test function. Then we have
$\Bigl|\int\limits_{X}\psi(\omega^{n}_{u_{j}}-\omega^{n}_{\phi_{j}})\Bigl|\leq\int\limits_{X}|\psi|\|\omega^{n}_{u_{j}}-\omega^{n}_{\phi_{j}}\|\leq\sup\limits_{X}|\psi|\int\limits_{X}\|\omega^{n}_{u_{j}}-\omega^{n}_{\phi_{j}}\|.$ (3.9)
Because of the hypothesis (ii) the right-hand side of (3.9) tends to 0 as $j\to\infty$ and the desired conclusion follows. ∎
Next, we give a following result which is an extension of Corollary 1 in [20] in Hermitian setting. Namely, we have the following result.
Proposition 3.10.
Let $(u_{j})\subset\mathcal{P}_{0}$ be a uniformly bounded sequence and $u\in\mathcal{P}_{0}$. Assume that $u_{j}\rightarrow u$ in $L^{1}(X,dV)$ and there exists a positive Radon measure $\mu$ on $X$ vanishing on pluripolar sets. If $\omega^{n}_{u_{j}}\leq\mu$ for all $j\geq 1$ then $\omega^{n}_{u_{j}}\rightarrow\omega^{n}_{u}$ weakly in $X$.
Proof.
It is enough to show that every subsequence of the $\omega^{n}_{u_{j}}$ has its subsequence which is weakly convergent to $\omega^{n}_{u}$. Therefore, without loss of generality, we consider $(u_{j})$ as a subsequence of the original sequence. Because $u_{j}\rightarrow u$ in $L^{1}(X,dV)$ then there exists a subsequence $(u_{j_{k}})\subset(u_{j})$ such that $u_{j_{k}}$ converges $dV$-a.e. to $u$. Then by Corollary 2.2 in [17] we can choose a subsequence of the $(u_{j_{k}})$ which we still denote it by $(u_{j_{k}})$ satisfying
$\lim\limits_{k\to\infty}\int\limits_{X}|u_{j_{k}}-u|d\mu=0.$
For an arbitrary $\delta>0$ consider
$\int\limits_{\{u\geq u_{j_{k}}+\delta\}}\omega^{n}_{u_{j_{k}}}$ $\leq\frac{1}{\delta}\int\limits_{\{u\geq u_{j_{k}}+\delta\}}|u_{j_{k}}-u|\omega^{n}_{u_{j_{k}}}$
$\leq\frac{1}{\delta}\int\limits_{X}|u_{j_{k}}-u|\omega^{n}_{u_{j_{k}}}\leq\frac{1}{\delta}\int\limits_{X}|u_{j_{k}}-u|d\mu\rightarrow 0,$
as $k\to\infty$. By using Theorem 3.3 we infer that $\omega^{n}_{u_{j_{k}}}\rightarrow\omega^{n}_{u}$ weakly in $X$. The proof is complete. ∎
4 Applications
In this section by relying on the main result in [17] we give a result about solving Monge-Ampère equation on compact Hermitian manifolds if there exists a smooth subsolution. Note that, for the case of a compact Hermitian manifold with boundary and with given boundary conditions, if we have the assumption about the existence of a subsolution, then the existence of a solution to the Dirichlet problem for the complex Monge-Ampère equation has been recently established in [16]. Now we state and prove the following result.
Theorem 4.1.
Let $(X,\omega)$ be a compact Hermitian manifold of complex dimension $n$, with smooth volume form $dV=\omega^{n}$ as in Section 3. Suppose that $\mu$ is a positive Borel measure with $\mu(X)=\int\limits_{X}\omega^{n}=\int\limits_{X}dV$ and there exists a function $v\in PSH(\omega)\cap C^{\infty}(X)$ such that $\mu\leq A\omega_{v}^{n}$ for some constant $A>0$. Then, there exists a function $u\in PSH(\omega)\cap L^{\infty}(X)$ and a constant $c>0$ such that $\omega^{n}_{u}=c\mu$.
Proof.
Using Lebesgue-Radon-Nikodym theorem we can write $\mu=f\omega^{n}_{v}$, $f\in L^{1}(X,\omega^{n}_{v}),0\leq f\leq A$. Next, we choose a sequence $(f_{j})\in L^{1}(X,\omega^{n}_{v})\cap C^{\infty}(X)$ such that $0<f_{j}\leq 2A$ for all $j\geq 1$ and $\int\limits_{X}|f_{j}-f|\omega^{n}_{v}\rightarrow 0$ as $j\to\infty$. Since the measure $f_{j}\omega^{n}_{v}$ is smooth, by a well-known result of [18], there exists a sequence $(u_{j})\in PSH(\omega)\cap C^{\infty}(X),\sup\limits_{X}u_{j}=0$ and $B_{j}>0$ such that
$\omega^{n}_{u_{j}}=B_{j}f_{j}\omega^{n}_{v}.$ (4.1)
According to Lemma 5.9 in [15], we infer that $0<B_{j}\leq B$ with a uniform $B>0.$ Therefore, it follows from Corollary 5.6 in [15] that we have $C\leq u_{j}\leq 0$ with uniform $C>0.$ By passing to a subsequence we may assume that $u_{j}\to u$ in $L^{1}(X,dV)$, $u\in PSH(\omega)$, $C\leq u\leq 0$ and $B_{j}\to c\geq 0.$ We have
$\omega^{n}_{u_{j}}=B_{j}f_{j}\omega^{n}_{v}\leq 2AB\omega^{n}_{v}.$ (4.2)
From (4.2) and utilizing the main result in [17], there is a subsequence of $\{u_{j}\},$ for simplicity still denoted by $\{u_{j}\}$ such that $\omega^{n}_{u_{j}}\to\omega^{n}_{u}$ as $j\to\infty.$ By Equation (4.1) we infer that
$\omega^{n}_{u}=cf\omega^{n}_{v}=c\mu.$
Since $u$ is bounded, by Remark 5.7 in [15] it follows that $c>0$. The proof of theorem is complete. ∎
Data Availibility Statement
The manuscript has no associated data.
Ethics declarations
Conflict of interest
The authors declare that no conflict of interest in this paper is reported.